a)calcualte the value of the discriminant of x^2+4x+4=0
b)By examining the sign of the discriminant in part a, how many x-intercepts would the graph of y=x^2+4x+4 have? why?

Answer by stanbon(75887)

a)calcualte the value of the discriminant of x^2+4x+4=0
b^2-4ac = 4^2-4*1*4 = 16-16 = 0
-------------------------------------
b)By examining the sign of the discriminant in part a, how many x-intercepts would the graph of y=x^2+4x+4 have?
It would touch the x-axis at x = -2 but would not pass through the axis.
-------------
why?
Since the discriminant is 0,
x = -b/2a = -4/(2*1) = -2 with multiplicity two.
Cheers,
Stan H.